Earley Corporation issued perpetual preferred stock with a 8% annual dividend. The stock currently yields 10%, and its par value is $100. What is the stock's value? Round your answer to two decimal places. $_____

Suppose interest rates rise and pull the preferred stock's yield up to 14%. What is its new market value? Round your answer to two decimal places. $_____
Answer #1

Value of preferred stock = Annual dividend / Annual yield = ($100 * 0.08) / 0.10 = $8/0.10 = $80 per share

New market value of preferred stock = $8 / 0.14 = $57.14 per share